a father is 42 years old. his children are 6,8 and 12 years old respectively. how old will the father be when his age equals the sum of his three children's ages ?
Answer by Alan3354(69443)

a father is 42 years old. his children are 6,8 and 12 years old respectively. how old will the father be when his age equals the sum of his three children's ages ?
--------
respectively?
---
The sum of their ages is 26
The sum increases 3 years per year.
---
26 + 3y = 42 + y
2y = 16
y = 8 years
----
--> 14 + 16 + 20 = 42 + 8